### v1

# YITP-25-129, KOBE-COSMO-25-16

YITP-25-129, KOBE-COSMO-25-16
Quantum nature of gravitational waves from binary black holes
Sugumi Kanno∗♮, Jiro Soda, and Akira Taniguchi
     Department of Physics, Kyushu University, Fukuoka 819-0395, Japan [5pt]       Center for Gravitational Physics and Quantum Information, [2pt]       Yukawa Institute for Theoretical Physics, Kyoto University, [2pt]       Department of Physics, Kobe University, Kobe 657-8501, Japan
Quantum mechanics is the fundamental framework of nature, and gravitational waves from binary black holes should likewise be analyzed quantum mechanically. It is commonly assumed that their classical description corresponds to a coherent state, so any deviation would signal genuinely quantum nature of gravity. We show that the coherent-state description reproduces classical gravitational waves at leading order, while next-order effects generate squeezed states of gravitons. For GW150914, we estimate the squeezing parameter to be $\sim 10^{-3}$. Detection of such squeezing with LIGO, Virgo, or KAGRA would provide direct evidence for the quantization of gravity.
1 Introductioin
On macroscopic scales, gravity plays a central role. One of the most remarkable predictions of general relativity is the existence of fluctuations in spacetime, known as gravitational waves. In 2015, gravitational waves from a binary black hole merger were directly detected for the first time [1].
It is widely accepted that quantum theory provides the fundamental framework governing all physical phenomena at any scale in the universe. Schrödinger’s cat has become an icon of this view. Upon quantization of gravitational waves, a new particle — the graviton — is predicted. Dyson discussed the detectability of a single graviton and concluded that such detection would be impossible in practice [2]. However, it has been pointed out that squeezed states may allow for an indirect probe of gravitons [3, 4, 5, 6, 7, 8] (see also a review [9]). Thus, it is worthwhile to explore the role of quantum states of gravitons on macroscopic scales.
From the perspective of the quantum state of gravitons, a key question is how classical gravitational waves should be described within quantum theory. By analogy with quantum optics, it has been implicitly assumed that classical gravitational waves correspond to a coherent state [10]. More recently, this idea has been formalized as the coherent state hypothesis [11, 12, 13].
In quantum optics, the state of photons is not necessarily a coherent state but can also be a squeezed state, generated through the nonlinear response of a medium. In this sense, photons exhibit distinctly quantum behavior. By analogy, one may expect similar phenomena for gravitons. In particular, squeezing of graviton states can occur in the presence of strong gravitational fields. A well-known example is the squeezed state of primordial gravitational waves generated during inflation [14, 15]. Another is the squeezed state associated with Hawking radiation from black holes [16]. Strong gravitational fields are also present in binary systems that emit gravitational waves. Curiously, however, gravitational waves from binary black holes have scarcely been discussed in the context of quantum theory. This omission is often attributed to the macroscopic nature of the system, where quantum effects are usually assumed to be negligible. Yet, such an assumption may reflect prejudice rather than necessity.
For progress in unifying quantum theory and gravity, it is crucial to reveal the non-classical aspects of gravity or, ultimately, to detect gravitons. Recently, the role of nonlinear effects in gravitational waves has been investigated within the framework of quantum theory [17, 18]. In addition, squeezed graviton states arising from superradiant axions have been discussed [19, 20]. It should be emphasized, however, that all gravitational waves detected so far originate exclusively from binary black holes. Hence, it is important to investigate the quantum nature of gravitational waves emitted by such systems. To this end, we first formulate a coherent state description of gravitational waves. We then derive a formula for the squeezing parameters. Furthermore, we estimate the squeezing parameter for the event GW150914 and discuss the prospects for detecting the quantum nature of gravitational waves, namely gravitons.
The organization of the paper is as follows. In Section 2, we present the basic framework for generating the quantum state of gravitational waves from binary black holes, and in particular derive the interaction Hamiltonian between a binary system and gravity. In Section 3, we formulate a coherent-state description of gravitational waves from binary black holes, and show that conventional results of quantum theory are successfully reproduced. In Section 4, we derive a formula for the squeezing parameters of the observable graviton state, which quantifies the deviation from a coherent state. As an application, we evaluate the squeezing parameter for the event GW150914 and obtain the intriguing value of $10^{-3}$. This result suggests the possibility of probing the quantum nature of gravitational waves, namely the existence of gravitons. The final section is devoted to conclusions. Throughout this work, we adopt natural units with $c=\hbar=1$.
2 Basic framework
In the case of photons, a classical current generates a coherent state of photons as
$\exp[-i\int dtd^{3}x{\bf j}(t)\cdot{\bf\hat{A}}(x^{i}(t),t)]\ ,$ (2.1)
where ${\bf j}(t)$ is a classical current and $^{\bf\hat{A}}$ is the operator evaluated along the trajectory. Since the interaction is linear, only a coherent state is produced.
By contrast, in the case of gravity, a classical source induces the coupling
$\exp[-i\int dtd^{3}x\left\{T_{ij}(t)\hat{h}_{ij}(x^{i}(t),t)+\Lambda_{ijkl}(t)\hat{h}_{ij}(x^{i}(t),t)\hat{h}_{kl}(x^{i}(t),t)+\cdots\right\}]\ ,$ (2.2)
where $\Lambda_{ijkl}(t)$ is a classical tensor determined by the interaction between matter and gravity. The linear term in $\hat{h}_{ij}$ generates a coherent state, while the quadratic term generates a squeezed state.
We consider a binary system with component masses $m_{1}$ and $m_{2}$. The motion of the black holes is described by the geodesics $\zeta_{1},\zeta_{2}$, respectively. For simplicity, we assume that the orbital trajectories $x_{1}^{i}(t)$ and $x_{2}^{i}(t)$ are given, neglecting the back-reaction due to gravitational wave emission.
The interaction of the corresponding energy-momentum tensor with the graviton field produces a coherent state of gravitons, which describes the classical gravitational waves. By including higher-order interactions, one can go beyond the coherent-state description. In this section, we derive the interaction Hamiltonian that governs the quantum state of gravitons emitted by the binary black holes.
The total action consists of the Einstein–Hilbert action together with the geodesic actions of the two particles with masses $m_{1}$ and $m_{2}$
$S\hskip-5.0pt$ $=$ $\hskip-5.0ptS_{\rm EH}+S_{1}+S_{2}=\frac{M_{\rm p}^{2}}{2}\int d^{4}x\sqrt{-g}R-m_{1}\int_{\zeta_{1}}d\tau-m_{2}\int_{\zeta_{2}}d\tau\ ,$ (2.3)
where $M_{\rm p}=1/\sqrt{8\pi G}$ is the Plank mass, $R$ is the Ricci scalar, and $\tau$ is the proper time.
We consider gravitons in the Minkowski space as tensor-mode perturbation of the spatial metric
$-d\tau^{2}=ds^{2}=-dt^{2}+\left(\delta_{ij}+h_{ij}\right)dx^{i}dx^{j}\ ,$ (2.4)
where $\delta_{ij}$ is the Kronecker delta and $h_{ij}$ is the metric perturbation, subject to the transverse-traceless conditions $h^{i}{}_{i}=h^{i}{}_{j,i}=0$. The indices $(i,j)$ run from 1 to 3, corresponding to $(x,y,z)$. Substituting the metric (2.4) into the action (2.3) and expanding to second order in $h_{ij}$, we obtain
$S_{\rm EH}=\frac{M_{\rm p}^{2}}{8}\int d^{4}x\left(\dot{h}_{ij}\dot{h}_{ij}-h_{ij,k}h_{ij,k}\right)\ ,$ (2.5)
From this quadratic action, we identify the canonical variable as $\psi_{ij}\equiv h_{ij}M_{\rm p}/2$ with the conjugate momentum given by $\dot{\psi}_{ij}$ .
We quantize the gravitational waves by imposing the canonical commutation relations,
$\left[\psi_{ij}(t,\bm{x}),\dot{\psi}^{k\ell}(t,\bm{x}^{\prime})\right]\hskip-5.0pt$ $=$ $\hskip-5.0pt\frac{i}{2}\left(P_{i}{}^{k}P_{j}{}^{\ell}+P_{i}{}^{\ell}P_{j}{}^{k}-P_{ij}P^{k\ell}\,\right)\delta\left(\bm{x}-\bm{x}^{\prime}\right)\,$, (2.6)
where the transverse projection tensor is defined as
$P_{ij}=\delta_{ij}-\frac{\partial_{i}\partial_{j}}{\nabla^{2}}~.$ (2.7)
Thus, the gravitational field $h_{ij}$ can be expanded as follows
$h_{ij}(t,\bm{x})=\frac{2}{M_{\rm p}}\sum_{P=+,\times}\int\frac{d^{3}\bm{k}}{(2\pi)^{3/2}}\left[\frac{e^{-i\omega_{\bm{k}}t}}{\sqrt{2\omega_{\bm{k}}}}e^{(P)}_{ij}(\bm{k})a^{(P)}(\bm{k})+\frac{e^{i\omega t}}{\sqrt{2\omega_{k}}}e^{(P)}_{ij}(-\bm{k})a^{(P){\dagger}}(-\bm{k})\right]e^{i\bm{k}\cdot\bm{x}}\,$, (2.8)
where $e^{(P)}_{ij}(\bm{k})\,,(P=+,\times)$ are the polarization tensors, normalized as $e^{(P)}_{ij}(\bm{k})e^{(Q)}_{ij}(\bm{k})=\delta^{PQ}$. The anihilation and creation operators satisfy the commutation relation
$\left[a^{(P)}(\bm{k}),a^{(Q){\dagger}}(\bm{k}^{\prime})\right]\hskip-5.0pt$ $=$ $\hskip-5.0pt\delta({\bm{k}}-{\bm{k}}^{\prime})\delta^{PQ}\ .$ (2.9)
The action for the geodesics motion of the particles $\gamma_{N},(N=1,2)$ contains the interaction terms of the form
$S_{\rm int}$ $=$ $-\sum_{N=1,2}m_{N}\int_{\zeta_{N}}\sqrt{dt^{2}-\delta_{ij}dx^{i}dx^{j}-h_{ij}(t,\bar{\bm{x}}_{N}(t))dx^{i}dx^{j}}\,$, (2.10)
$=$ $-\sum_{N=1,2}m_{N}\int_{\zeta_{N}}dt~\frac{1}{\gamma_{N}}\sqrt{1-\gamma_{N}^{2}h_{ij}(t,\bar{\bm{x}}_{N}(t))v_{N}^{i}v_{N}^{j}}\ .$
Here, $\bar{\bm{x}}_{N}(t)$ denotes the trajectory of the $N$-th particle, $v_{N}^{i}=d\bar{x}_{N}^{i}/dt$ is its velocity with $v_{N}^{2}=v_{N}^{i}v_{N}^{i}$, and $\gamma_{N}=1/\sqrt{1-v_{N}^{2}}$ is the Lorentz factor. Therefore, after performing the Legendre transformation, the interaction Hamiltonian up to second order in $h_{ij}$ takes the form
$H_{\rm int}(t,\bar{\bm{x}})\hskip-5.0pt$ $=$ $\hskip-5.0pt\sum_{N=1,2}\left[\frac{\gamma_{N}^{3}m_{N}}{2}h_{ij}(t,\bar{\bm{x}}_{N}(t))v_{N}^{i}v_{N}^{j}\right$. (2.11)
$\left.\hskip 56.9055pt+\frac{3}{8}\gamma_{N}^{5}m_{N}h_{ij}(t,\bar{\bm{x}}_{N}(t))\ h_{lm}(t,\bar{\bm{x}}_{N}(t))\ v_{N}^{i}v_{N}^{j}v_{N}^{l}v_{N}^{m}\right]\,.$
We work in the interaction picture, where the time evolution operator $\hat{U}(t,\bar{\bm{x}})$ for the graviton quantum state is governed by the interaction Hamiltonian $\hat{H}_{\rm int}$:
$\hat{U}(t,\bar{\bm{x}})\hskip-5.0pt$ $=$ $\hskip-5.0pt\mathcal{T}\left[\exp\left(-i\int^{t}dt^{\prime}\hat{H}_{\rm int}(t^{\prime})\right)\right]$ (2.12)
where $\mathcal{T}$ denotes time ordering.
3 Coherent state description of gravitatinal waves
We assume circular orbital motion. Taking the center of mass as the origin, the tragectories can be written as
$\begin{array}[]{lll}x_{1}=\displaystyle\frac{m_{2}}{M}a\cos(\Omega t)\ ,&y_{1}=\displaystyle\frac{m_{2}}{M}a\sin(\Omega t)\ ,&z_{1}=0\ ,\\[10.0pt] x_{2}=\displaystyle\frac{m_{1}}{M}a\cos(\Omega t+\pi)\ ,&y_{2}=\displaystyle\frac{m_{1}}{M}a\sin(\Omega t+\pi)\ ,&z_{2}=0\ ,\end{array}$ (3.3)
where $M=m_{1}+m_{2}$ is the total mass and $a$ is the orbital separation. Introducing the reduced mass $\mu=(m_{1}m_{2})/M$, we obtain the useful relation $m_{1}x_{1}^{2}+m_{2}x_{2}^{2}=\mu a^{2}$. This shows that the reduced mass $\mu$ effectively characterizes the binary system’s quadrupole moment, which plays a central role in the generation of gravitational waves.
The time evolution operator describes the quantum state of gravitons produced by the binary black holes. As discussed in Section 2, retaining only the linear term in $h_{ij}$ in the interaction Hamiltonian yields an operator that generates a coherent state
$\hat{U}(t,\bar{\bm{x}}_{N})\hskip-5.0pt$ $=$ $\hskip-5.0pt\mathcal{T}\left[\exp\left(-i\int^{t}dt^{\prime}\hat{H}_{\rm int}(t^{\prime})\right)\right]$ (3.4)
$=$ $\hskip-5.0pt\exp\left[-i\frac{2}{M_{\rm p}}\sum_{N=1,2}\frac{\gamma_{N}^{3}m_{N}}{2}\sum_{P=+,\times}\int^{t}dt^{\prime}\int\frac{d^{3}\bm{k}}{(2\pi)^{3/2}}\right$.
$\hskip 28.45274pt\left.\times\left(\frac{e^{-i\omega_{\bm{k}}t^{\prime}}}{\sqrt{2\omega_{\bm{k}}}}e^{(P)}_{ij}(\bm{k})a(\bm{k})+\frac{e^{i\omega_{\bm{k}}t^{\prime}}}{\sqrt{2\omega_{\bm{k}}}}e^{(P)}_{ij}(-\bm{k})a^{\dagger}(-\bm{k})\right)e^{i\bm{k}\cdot\bar{\bm{x}}_{N}}v_{N}^{i}v_{N}^{j}\right]\ .$
Comparing the above expression with the definition of the displacement operator
$\hat{D}(\alpha)=\prod_{P}\exp\left[\int d^{3}{\bm{k}}\left(\alpha^{(P)}({\bm{k}})a^{(P){\dagger}}({\bm{k}})-\alpha^{(P)*}({\bm{k}})a^{(P)}({\bm{k}})\right)\right]\ ,$ (3.5)
we identify the coherent state parameter as
$\alpha^{(P)}(\bm{k})\hskip-5.0pt$ $=$ $\hskip-5.0pt-\frac{i}{(2\pi)^{3/2}}\sum_{N=1,2}\int^{t}dt^{\prime}\frac{\gamma_{N}^{3}m_{N}}{M_{\rm p}}\frac{e^{i\omega_{\bm{k}}t^{\prime}}}{\sqrt{2\omega_{\bm{k}}}}e^{(P)}_{ij}(\bm{k})v_{N}^{i}v_{N}^{j}e^{-i\bm{k}\cdot\bar{\bm{x}}_{N}}\ .$ (3.6)
This parameter encodes the classical orbital dynamics of the binary system into the quantum coherent state of gravitons, thereby providing the bridge between the classical gravitational wave signal and its quantum description.
Choosing the $z$-axis in $\bm{k}$} space to align with the position vector $\bm{x}$}, we parametrize the wave vector as $\bm{k}=k(\sin\theta\cos\varphi,\sin\theta\sin\varphi,\cos\theta)$. The coherent state parameter evaluate to
$\alpha^{(+)}(\bm{k})\hskip-5.0pt$ $=$ $\hskip-5.0pt\frac{i}{(2\pi)^{3/2}}\frac{\mu(a\Omega)^{2}}{\sqrt{2}M_{\rm p}}\int^{t}dt^{\prime}\frac{e^{i\omega_{\bm{k}}t^{\prime}}}{\sqrt{2\omega_{\bm{k}}}}\left(\frac{\sin^{2}\theta}{2}+\frac{1+\cos^{2}\theta}{2}\cos(2\Omega t^{\prime}-2\varphi)\right)$ (3.7)
$\hskip 99.58464pt\times\left[\gamma^{3}_{1}\left(\frac{m_{2}}{M}\right)e^{-i(k_{x}x_{1}+k_{y}y_{1})}+\gamma^{3}_{2}\left(\frac{m_{1}}{M}\right)e^{-i(k_{x}x_{2}+k_{y}y_{2})}\right]\ ,$
$\alpha^{(\times)}(\bm{k})\hskip-5.0pt$ $=$ $\hskip-5.0pt\frac{i}{(2\pi)^{3/2}}\frac{\mu(a\Omega)^{2}}{\sqrt{2}M_{\rm p}}\int^{t}dt^{\prime}\frac{e^{i\omega_{\bm{k}}t^{\prime}}}{\sqrt{2\omega_{\bm{k}}}}\cos\theta\sin(2\Omega t^{\prime}-2\varphi)$ (3.8)
$\hskip 99.58464pt\times\left[\gamma^{3}_{1}\left(\frac{m_{2}}{M}\right)e^{-i(k_{x}x_{1}+k_{y}y_{1})}+\gamma^{3}_{2}\left(\frac{m_{1}}{M}\right)e^{-i(k_{x}x_{2}+k_{y}y_{2})}\right]\ ,$
where
$k_{x}x_{1}+k_{y}y_{1}\hskip-5.0pt$ $=$ $\hskip-5.0pt\frac{m_{2}}{M}ak(\sin\theta\cos\varphi\cos(\Omega t^{\prime})+\sin\theta\sin\varphi\sin(\Omega t^{\prime}))\ ,$ (3.9)
$k_{x}x_{2}+k_{y}y_{2}\hskip-5.0pt$ $=$ $\hskip-5.0pt-\frac{m_{1}}{M}ak(\sin\theta\cos\varphi\cos(\Omega t^{\prime})+\sin\theta\sin\varphi\sin(\Omega t^{\prime}))\ .$ (3.10)
We also used the explicit form of polarization tensors
$e^{(+)}_{ij}\hskip-5.0pt$ $=$ $\hskip-5.0pt\frac{1}{\sqrt{2}}\begin{pmatrix}\cos^{2}\theta\cos^{2}\varphi-\sin^{2}\varphi&(1+\cos^{2}\theta)\sin\varphi\cos\varphi&-\displaystyle\frac{1}{2}\sin 2\theta\cos\varphi\\[12.0pt] (1+\cos^{2}\theta)\sin\varphi\cos\varphi&\cos^{2}\theta\sin^{2}\varphi-\cos^{2}\varphi&-\displaystyle\frac{1}{2}\sin 2\theta\sin\varphi\\[12.0pt] -\displaystyle\frac{1}{2}\sin 2\theta\cos\varphi&-\displaystyle\frac{1}{2}\sin 2\theta\sin\varphi&\sin^{2}\theta\end{pmatrix}$, (3.11)
$e^{(\times)}_{ij}\hskip-5.0pt$ $=$ $\hskip-5.0pt\frac{1}{\sqrt{2}}\begin{pmatrix}-\cos\theta\sin 2\varphi&\cos\theta\cos 2\varphi&\sin\theta\sin\varphi\\[12.0pt] \cos\theta\cos 2\varphi&\cos\theta\sin 2\varphi&-\sin\theta\cos\varphi\\[12.0pt] \sin\theta\sin\varphi&-\sin\theta\cos\varphi&0\end{pmatrix}\$. (3.12)
[…]
The expectation value of the metric operator in the coherent state is
[…]
Substituting the coherent parameters into the above expression, we obtain
[…]
$\times\left\{i\left[\gamma^{3}_{1}\left(\frac{m_{2}}{M}\right)e^{-i(k_{x}x_{1}+k_{y}y_{1})}+\gamma^{3}_{2}\left(\frac{m_{1}}{M}\right)e^{-i(k_{x}x_{2}+k_{y}y_{2})}\right]e^{-ikr\cos\theta+i\omega_{k}t}e^{i\omega_{k}t^{\prime}}\right$.
$\hskip-42.67912pt\left.\times\left[e_{ij}^{(+)}\left(\frac{\sin^{2}\theta}{2}+\frac{1+\cos^{2}\theta}{2}\cos(2\Omega t^{\prime}-2\varphi)\right)+e_{ij}^{(\times)}\cos\theta\sin\left(2\Omega t^{\prime}-2\varphi\right)\right]+{\rm c.c.}\right\}\ ,\qquad$
where $r=|\bm{x}|$ and the integrals can be evaluated approximately. For $r\Omega\gg 1$, we obtain
[…]
Interestingly, this expression contains not only the expected outgoing waves but also ingoing waves. In addition, when compared with the standard quadrupole formula, the amplitude differs by a factor of two. If we add the contributions from both the outgoing and ingoing waves, the correct amplitude is recovered. It is therefore important to clarify the origin of this apparent discrepancy.
The presence of the ingoing component arises because the coherent-state expectation value incorporates all of the directions of wavenumber vectors of the field, corresponding to advanced and retarded solutions of the wave equation. In the standard classical treatment, one imposes retarded boundary conditions to eliminate the ingoing contribution. The discrepancy in amplitude therefore reflects the fact that the coherent-state construction, taken at face value, does not yet enforce the choice of purely retarded (outgoing) solutions.
Now, we can estimate the amplitude of gravitational waves at a distance $r$ from the binary system. As a reference, let us use the parameters of GW150914 [1]. The component Black hole masses are $36~M_{\odot}$ and $29~M_{\odot}$, giving the reduced mass of $\mu\sim 16~M_{\odot}$. The luminosity distance to the source is $410~{\rm Mpc}$. We consider the orbit of the black holes at the innermost stable circular orbit (ISCO), where the orbital velocity is $a\Omega=1/\sqrt{6}\sim 0.41$. In this case, the Lorentz factors are $\gamma_{1}=1.02$ and $\gamma_{2}=1.03$. Substituting these values, we obtain
[…]
This estimate is consistent with the amplitude of the gravitational waves detected in GW150914.
4 Nonclassicality of gravitational waves
At the next order of the perturbative expansion in Section 3, the time evolution operator aquires terms that generates a squeezed state. As discussed in Section 2, such squeezing arises from the term quadratic interaction term proportional to $h_{ij}^{2}$. The corresponding operator takes the form
$\hskip-14.22636ptU(t,\bar{\bm{x}})\hskip-5.0pt$ $=$ $\hskip-5.0pt\exp\left[-i\sum_{N=1,2}\int^{t}dt^{\prime}\frac{3\gamma_{N}^{5}m_{N}}{2M_{\rm p}^{2}}\sum_{P,Q}\int\frac{d^{3}{\bm{k}}}{(2\pi)^{3/2}}\int\frac{d^{3}{\bm{k}}^{\prime}}{(2\pi)^{3/2}}\right$. (4.1)
$\hskip-14.22636pt\left.\times\left(\frac{e^{-i(\omega_{k}+\omega_{k^{\prime}})t^{\prime}}}{2\sqrt{\omega_{k}\omega_{k^{\prime}}}}e^{(P)}_{ij}({\bm{k}})e^{(Q)}_{lm}({\bm{k}}^{\prime})v_{N}^{i}v_{N}^{j}v_{N}^{l}v_{N}^{m}a^{(P)}({\bm{k}})a^{(Q)}({\bm{k}}^{\prime})e^{i\bm{k}\cdot\bar{\bm{x}}_{N}}e^{i\bm{k}^{\prime}\cdot\bar{\bm{x}}_{N}}+{\rm h.c.}\right)\right]\ .$
The terms proportional to $aa^{\dagger}$ or $a^{\dagger}a$ correspond only to phase rotations and do not contribute to the squeezing amplitude. We therefore neglect them.
Comparing this operator with the definition of the squeeze operator,
$\hat{S}(\beta)=\prod_{P}\exp\left[\int d^{3}{\bm{k}}\int d^{3}{\bm{k}}^{\prime}\left(\beta^{(P)}_{kk^{\prime}}a^{(P){\dagger}}({\bm{k}})a^{(P){\dagger}}({\bm{k}}^{\prime})-\beta_{kk^{\prime}}^{(P)*}a^{(P)}({\bm{k}})a^{(P)}({\bm{k}}^{\prime})\right)\right]\ ,$ (4.2)
we can identify the squeezing parameter as
$\hskip-28.45274pt\beta^{(P)}_{kk^{\prime}}\hskip-5.0pt$ $=$ $\hskip-5.0pt-\frac{3i}{(2\pi)^{3}2M_{\rm p}^{2}}\sum_{N=1,2}\gamma_{N}^{5}m_{N}\int^{t}dt^{\prime}\frac{e^{i(\omega_{k}+\omega_{k^{\prime}})t^{\prime}}}{2\sqrt{\omega_{k}\omega_{k^{\prime}}}}e^{(P)}_{ij}({\bm{k}})e^{(P)}_{lm}({\bm{k}}^{\prime})v_{N}^{i}v_{N}^{j}v_{N}^{l}v_{N}^{m}e^{-i\bm{k}\cdot\bar{\bm{x}}_{N}}e^{-i\bm{k}^{\prime}\cdot\bar{\bm{x}}_{N}}\ .$ (4.3)
Here we have also noted that cross terms involving ++ and $\times$ polarizations vanish after averaging over wave numbers. By using Eqs. (3.3),(3.11) and (3.12), the squeezing parameter can be evalulated explicitly as
$\beta^{(+)}_{kk^{\prime}}\hskip-5.0pt$ $=$ $\hskip-5.0pt-\frac{3i}{64\pi^{3}M_{\rm p}^{2}}\mu(a\Omega)^{4}$ (4.4)
$\hskip-28.45274pt\times\int^{t}dt^{\prime}\left[\gamma_{1}^{5}\left(\frac{m_{2}}{M}\right)^{3}e^{-i(k_{x}+k^{\prime}_{x})x_{1}-i(k_{y}+k^{\prime}_{y})y_{1}}+\gamma_{2}^{5}\left(\frac{m_{1}}{M}\right)^{3}e^{-i(k_{x}+k^{\prime}_{x})x_{2}-i(k_{y}+k^{\prime}_{y})y_{2}}\right]$
$\hskip-42.67912pt\times\frac{e^{i(\omega_{k}+\omega_{k^{\prime}})t^{\prime}}}{\sqrt{\omega_{k}\omega_{k^{\prime}}}}\left[\frac{\sin^{2}\theta}{2}+\frac{1+\cos^{2}\theta}{2}\cos(2\Omega t^{\prime}-2\varphi)\right]\left[\frac{\sin^{2}\theta^{\prime}}{2}+\frac{1+\cos^{2}\theta^{\prime}}{2}\cos(2\Omega t^{\prime}-2\varphi^{\prime})\right]\$
and
$\beta^{(\times)}_{kk^{\prime}}\hskip-5.0pt$ $=$ $\hskip-5.0pt-\frac{3i}{64\pi^{3}M_{\rm p}^{2}}\mu(a\Omega)^{4}$ (4.5)
$\times\int^{t}dt^{\prime}\left[\gamma_{1}^{5}\left(\frac{m_{2}}{M}\right)^{3}e^{-i(k_{x}+k^{\prime}_{x})x_{1}-i(k_{y}+k^{\prime}_{y})y_{1}}+\gamma_{2}^{5}\left(\frac{m_{1}}{M}\right)^{3}e^{-i(k_{x}+k^{\prime}_{x})x_{2}-i(k_{y}+k^{\prime}_{y})y_{2}}\right]$
$\hskip 56.9055pt\times\frac{e^{i(\omega_{k}+\omega_{k^{\prime}})t^{\prime}}}{\sqrt{\omega_{k}\omega_{k^{\prime}}}}\cos\theta\sin(2\Omega t^{\prime}-2\varphi)\cos\theta^{\prime}\sin(2\Omega t^{\prime}-2\varphi^{\prime})\ .$
Constraints on the squeeing parameters have been discussed in the literature [21], based on LIGO data [22, 22].
We now evaluate the squeezing parameter at a location far from the binary system. The expressions derived above correspond to the values at the source. In this paper, we assume that no decoherence processes act on the gravitational waves during their propagation to the observer. Under this assumption, the value of the squeezing parameter at the detector is identical to that at the source. From Eqs. (4.4) and (4.5), we see that the squeezing parameter attains its largest value is obtained in the case where ${\bm{k}}=-{\bm{k}^{\prime}}$ and $\omega_{k}=\omega_{k^{\prime}}=k=2\Omega$. The squeezing parameter at the location of the binary system can be estimated as
$\zeta\simeq\frac{4\pi}{3}(2\Omega)^{3}|\beta|\simeq\frac{1}{8\pi M_{\rm p}^{2}}\mu(a\Omega)^{4}f\ .$ (4.6)
In the second expression, we have used the relation $\omega=2\pi f$. As in Section 3, we adopt the parameters of GW150914 [1]. In this case, we obtain
$\zeta\simeq 2\times 10^{-3}\left(\frac{\mu}{16~M_{\odot}}\right)\left(\frac{a\Omega}{0.41}\right)^{4}\left(\frac{f}{68~{\rm Hz}}\right)\ .$ (4.7)
Thus, the quantum state of gravitational waves from GW150914 is characterized by a squeezing parameter of order $10^{-3}$. While this estimate is based on the ISCO, the squeezing is expected to be even stronger just before the merger. For example, in GW150914 the frequency of the gravitational wave reached up to $150~{\rm Hz}$.
5 Conclusion
In this work, we investigated how astrophysical binary black holes generate a quantum state of gravitons. We modeled the binary system as a classical source. As is well known, when a classical source couples linearly to a quantum field, it produces a coherent state.
We confirmed that classical gravitational waves can be described within the coherent state framework. We then extended the analysis to the next-order effects and showed that binary black holes can also generate squeezed states. Furthermore, we estimated the degree of squeezing and found that, for GW150914, the squeezing parameter is of order $10^{-3}$. This result indicates a measurable deviation from a purely coherent state. Consequently, there may be a realistic opportunity to probe the nonclassicality of gravitational waves with current and future detectors such as LIGO, Virgo, and KAGRA.
Since we have obtained a squeezed coherent state, Hanbury–Brown–Twiss interferometry may be employed to probe the nonclassicality of the graviton state [23, 24, 25] with new technology [26, 27]. It would also be interesting to extend our analysis to other sources of gravitational waves. In this paper, we have demonstrated how a quantum state is generated by binary black holes under the assumption of the standard vacuum. However, if primordial gravitational waves generated during inflation are taken into account, an additional enhancement of squeezing is expected. Thus, by observing the squeezing of gravitational waves from binary black holes, one may obtain indirect information about the early universe. The detail will be reported separately [28].
Acknowledgments
S. K. was supported by the Japan Society for the Promotion of Science (JSPS) KAKENHI Grant Numbers JP22H01220, 24K21548 and MEXT KAKENHI Grant-in-Aid for Transformative Research Areas A “Extreme Universe” No. 24H00967. J. S. was in part supported by JSPS KAKENHI Grant Numbers JP23K22491, JP24K21548, JP25H02186. A. T. was supported by JSPS KAKENHI Grant Number JP25KJ1912.